OpenStudy (anonymous):

HOW TO FACTOR X TO THE SECOND POWER MINUS 9X PLUS 8?

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \bf x^2-9x+8=0 }\) \(\Huge\color{blue}{ \bf (x-8)(x-1)=0 }\) b/c (-8) * ( -1) = 8 AND -8 -1 =-9